OpenStudy (anonymous):

A bank pays 5% interest compounded annually. What principal will grow to $170,000 in 9 years? (Round your answer to two decimal places.)

OpenStudy (anonymous):

A=P(1+R/N)^NT P=A/(1+r/n)^nt P=170000/(1+.05/1)^(1*9)

OpenStudy (anonymous):

\[170000(1.05)^9\]

OpenStudy (anonymous):

i get $263,726 rounded

OpenStudy (anonymous):

$263,726 can't grow to $170,000